OpenStudy (anonymous):

Help with rational exponents: (x^1/4)^4/9

OpenStudy (anonymous):

Recall the following rule: \[\huge (x^m)^n=x^{mn}\]What do you think this means as it applies to your situation?

OpenStudy (anonymous):

Its \[x^ 1/9\]

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Wow, I just forgot to simplify... :) Thank you.

OpenStudy (anonymous):

\[x ^{\frac{1}{4}\times\frac{4}{9}}=x ^{\frac{4}{18}}=x ^{\frac{1}{9}}\]